# LIR: A Lightweight Baseline for Image Restoration

### v2

@@ Abstract
Recently, there have been significant advancements in Image Restoration based on CNN and transformer. However, the inherent characteristics of the Image Restoration task are often overlooked. Many works, instead, only focus on the basic block design and stack numerous such blocks to the model, leading to parameters redundant and computations unnecessary. Thus, the efficiency of the image restoration is hindered. In this paper, we propose a Lightweight Baseline for Image Restoration called LIR to efficiently reconstruct the image and remove degradations (blur, rain, noise, haze). First of all, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks, through a simple structural design. Then, to achieve lightweight, a Lightweight Adaptive Attention (LAA) Block is introduced depending on the inherent characteristics of the Image Restoration, which is mainly composed of proposed Adaptive Filters and Attention Blocks. LAA is capable of adaptively sharpening contours, removing degradation, and capturing global information in various Image Restoration scenes in a computation-friendly manner. Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art models with fewer parameters and computations in certain tasks. In addition, it is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic. The code is available at: https://github.com/Dongqi-Fan/LIR
Keywords:
@@
Image Restoration Lightweight Attention
[…]
[width=1]figures/introduction.eps
@@
Figure 1: Visualization experiment of the image restoration process. (a) The paradigm of the most modern networks that incorporate both local and global residual connections. (b) The output by Restormer still exhibits visible traces of rain. (c) The output of our LIR. Which is more plain and solves the problem of residual connections carrying degradation. (d) The visualization of the three levels of the encoder and decoder based on Restormer. The top is deblurring and bottom is deraining.
1 Introduction
@@
Image Restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, etc. In recent years, we have witnessed the great process of deep learning methods [28, 59, 52, 33, 39, 54, 67] in IR. However, many existing works only focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task. This approach often leads to parameter redundancy and computations unnecessary. Instead, our goal is an efficient and lightweight network design for various IR tasks.
@@
To address the issue of vanishing gradient and stabilize training, as shown in Fig. 1 (a), most modern networks usually incorporate local and global residual connections. However, we think that these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output. To validate our idea, we conduct visualization experiments on the deblurring and deraining tasks using Restormer [58], a transformer model with a U-Net structure. Fig. 1 illustrates the feature maps output by the encoder and decoder at three levels, as well as the input and output images of the Restormer, on the deblurring (at the top of Fig. 1 (d)) and deraining (at the bottom of Fig. 1 (d)) task. Notably, at the bottom of Fig. 1 (d), it is evident that each feature map and the final output of the network (Fig. 1 (b)) all contain degradation (rain). Although less pronounced in the deblurring task, a similar phenomenon can be observed at the top of Fig. 1 (d). To overcome this issue, we use an ingenious structural design (transpose branchs in the LIR and LAA Fig. 2) that effectively eliminates the degradation in the local and global residual connections. Fig. 1 (c) shows the restored image output by our LIR, which is more plain than the Restormer (Fig. 1 (b)).
@@
Moreover, we observe another serious issue in the visualization. In the visualization of the deblurring task (at the top of Fig. 1 (d)), the feature maps generated by encoder_level2 and decoder_level2 exhibit clear contours while the other feature maps exhibit blurred contours. This issue is even worse in the deraining (at the bottom of Fig. 1 (d)) task. All feature maps not only display very blurry contours but also contain a considerable amount of noise. Especially for the feature map generated by decoder_level1, it shows barely discernible contours, only with faint traces of rain. The more important is that the high-frequency information has been lost. These observations indicate that although many modern networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and fails to address the core issues of image restoration, thus resulting in poor visual quality. Based on this insight, we propose the Adaptive Filter (Fig. 3) (an component of the LAA) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
@@
Transformer-based models have demonstrated impressive performance in various tasks, both at high-level [12, 9, 31, 49] and low-level [6, 3, 54, 51, 27]. However, there exist different opinions. [10] argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs). In addition, [29, 47, 48] share a similar viewpoint that highlights the potential of MLPs beyond transformers. On the other hand, [50] suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy. Thus, excessive computations can be avoided. Furthermore, [16, 32, 8, 70] take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer. Satisfactory results are achieved and excessive computations are avoided. Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module (Fig. 4) (an component of the Attention Block) that aims to capture global information of the features and keeps parameters and computations friendly.
@@
In this paper, we propose a Lightweight Image Restoration network called LIR (Fig. 2) to efficiently remove degradation and adapt to various IR tasks. The key component of LIR is the novel Lightweight Adaptive Attention Block (LAA) consisting of several proposed Adaptive Filters and Attention Blocks. Futhermore, the Attention Block is formed by proposed Patch Attention module (Fig. 4), ResCABlock, and ResBlock (Details can be found in Fig. 2). Thanks to the Adaptive Filter and Attention Block, the LAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information. Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and LAA Block. This subtraction allows us to obtain the degradation, which is then expanded through the Transpose operation. After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection. Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art models in PSNR and produces better visual results that are more in line with the human aesthetic.
@@
  • Through a simple structural design, the problem that local and global residual connections bring degradation is addressed, which is overlooked by modern networks.
  • Given the problem of the loss of high-frequency information observed in the visualization experiments, we propose the Adaptive Filter (Fig. 3) as a solution to adaptively extract high-frequency information and enhance object contours.
  • To give LIR the ability to extract global information, we propose the Patch Attention module (Fig. 4), which is more computation-friendly than the self-attention of the transformer.
@@
2 Related Work
2.1 Image Restoration
In recent years, IR based on deep learning methods [13, 61, 55, 26, 63, 19] have shown promising results. The transformer architectures achieve adaptive spatial aggregation at the cost of heavy computation. To address this challenge, a Kernel Basis Attention (KBA) module has been proposed by KBNet [67] in the CNN. It adaptively aggregates spatial neighborhood information in a kernel-wise manner, thereby reducing computation. ADMS [38] proposes a Filter Attribution method based on Integral Gradient (FAIG) [55] to identify the contributions of the filters to remove specific degradations. By masking filters with minimal impact on degradation removal, the network’s efficiency in removing degradations is improved. IDR [61] proposes a learnable Principal Component Analysis (PCA) and views various IR tasks as multi-task learning methods to obtain network priors and remove degradations. MPRNet [59] introduces the Multi-Stage Progressive Restoration method, which is based on the U-Net architecture and follows a multi-input multi-output style. In addition, the contrastive learning paradigm is utilized by AirNet [24] to train the network to deal with different unknown degradations. Through the prompting of the contrastive learning paradigm, the latent features in the network gradually away from the degradation space and close to the clean space. Different from the above works, our emphasis is on lightweight and efficient design while ensuring good performance.
2.2 Vision Transformer for Image Restoration
IPT [3] is the first work that introduces ViT [12] to low-level tasks to achieve image super-resolution, denoising, and deraining. The computational complexity of the self-attention part in ViT is $O(N^{2}D)$, where $N=H$$\times$$W$ represents the length of the patch sequence and $D$ represents the dimension of the sequence. Although vision transformers are capable of capturing long-range dependencies, the self-attention component imposes a significant computing burden. The computational complexity of the self-attention increases quadratically with the input size. To mitigate this issue, TransWeather [53] introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$. Uformer [54] and StripFormer [51] proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size. To further reduce the computational complexity, [22] estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL [27] proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention. The computational complexities are reduced to $O(NlogND)$ and $O(NMD)$ respectively, where M represents the size of feature map. However, the computation required for self-attention remains unaffordable. To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings.
[…]
[width=]figures/model_2.eps
@@ 2.2 Vision Transformer for Image Restoration
Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Lightweight Adaptive Attention (LAA) Blocks. Thanks to the Adaptive Filter, Patch Attention, ResCABlock, and the ResBlock, the LAA Block is endowed with the ability to extract a variety of information efficiently.
3 Method
3.1 Lightweight Image Restoration
The overall structure of LIR is illustrated in Fig. 2, which primarily consists of Lightweight Adaptive Attention (LAA) Blocks. Both LIR and LAA Block adopt a down-up style to reduce computation and save parameters, that is downsizing the feature map in the beginning and upsizing at the end. To preserve low-frequency information and ensure training is stable, we add two global residual connections into LIR. One connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv.
As mentioned in Sec. 1, these residual connections can bring degradation. To address this, we subtract the output of the fourth LAA Block from the output of the first LAA Block to obtain the corresponding degradation. We then amplify the degradation through the Transpose operation and subtract it from the residual connection to obtain a clean connection. The same approach is used to remove degradation in the local residual connection within the LAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation. The reason we adopt this method is that the feature map of the shallow layer contains more degradation compared to a deep layer. Thus, this design allows the network to dynamically perceive this difference between shallow and deep layers during the training process and inference stage. Now, the problem that local and global residual connections bring degradation is addressed. Next, we illustrate how we can achieve lightweight and efficiency through the Adaptive Filter and Attention Block in detail.
[…]
[width=]figures/filters_2.eps
@@ 3.1 Lightweight Image Restoration
Figure 3: The comparison between our Adaptive Filter (left) and Edge-oriented Convolution Block (right). Both are reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
3.2 Adaptive Filter
@@
To achieve lightweight and efficiency, we first introduce the Adaptive Filter (at the left of Fig. 3), inspired by Edge-oriented Convolution Block (ECB) [66] (at the right of Fig. 3). The rich filters of different kinds allow the AF to enhance the object contour, remove degradation, and extract useful high-frequency information. Furthermore, the LAA block can be reparameterized as one $3\times 3$ and one high-pass filter after training to achieve lightweight.
@@
Although Adaptive Filter shares similarities with the ECB, it significantly outperforms ECB in three ways. Firstly, the ECB introduces more convolution operations during training leading to more training costs (i.e. parameters and computations). Instead, our Adaptive Filter only requires 10 training parameters resulting in smaller training costs. Secondly, the Adaptive Filter can be applied to various IR scenes and dynamically adjust the dependence on different filters by adjusting parameters $\alpha$. However, the ECB treats each branch equally and can not adapt to various IR scenes. Lastly, the Adaptive Filter offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
@@
Each filter in the Adaptive Filter has its own advantages and disadvantages. Thus, we use the learnable parameters $\alpha$ to adaptively maximize their advantages and minimize their disadvantages in different LAA blocks of the LIR. Specifically, in these filters, the Roberts filter is used to detect the pixel changes in the features by the differential method. It is effective in extracting prominent edges but fails to detect details. Where Roberts-x and Roberts-y represent the detection in the horizontal direction and vertical direction; The Sobel filter extracts edge information by approximating first-order differentials discretely. It has some resistance to noise and a good ability to detect details. However, the Sobel filter may cause edge thinning and false detection in scenes with a mass of noise; The Laplacian filter extracts edge information by approximating second-order differentials discretely. It can detect more detailed edges and textures than Sobel but is more sensitive to noise and may result in a bilateral effect on the edges. Where Laplacian-4 and Laplacian-8 represent 4 domains and 8 domains; The high-pass filter obtains high-frequency information by subtracting the features after Gaussian blur processing from the original features. This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the AF for stages where the features have scarce noise and a clear outline.
3.3 Attention Block

### v3

@@ Abstract.
Recently, there have been significant advancements in Image Restoration based on CNN and transformer. However, the inherent characteristics of the Image Restoration task are often overlooked in many works. They, instead, tend to focus on the basic block design and stack numerous such blocks to the model, leading to parameters redundant and computations unnecessary. Thus, the efficiency of the image restoration is hindered. In this paper, we propose a Lightweight Baseline network for Image Restoration called LIR to efficiently restore the image and remove degradations. First of all, through an ingenious structural design, LIR removes the degradations existing in the local and global residual connections that are ignored by modern networks. Then, a Lightweight Adaptive Attention (LAA) Block is introduced which is mainly composed of proposed Adaptive Filters and Attention Blocks. The proposed Adaptive Filter is used to adaptively extract high-frequency information and enhance object contours in various IR tasks, and Attention Block involves a novel Patch Attention module to approximate the self-attention part of the transformer. On the deraining task, our LIR achieves the state-of-the-art Structure Similarity Index Measure (SSIM) and comparable performance to state-of-the-art models on Peak Signal-to-Noise Ratio (PSNR). For denoising, dehazing, and deblurring tasks, LIR also achieves a comparable performance to state-of-the-art models with a parameter size of about 30%. In addition, it is worth noting that our LIR produces better visual results that are more in line with the human aesthetic. The code is available at: https://github.com/Dongqi-Fan/LIR
@@
Image Restoration; Patch Attention; Adaptive Filter; Residual Connection; Lightweight
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computational photographyccs: Computing methodologies Image processing
@@
Figure 1. Visualization experiment on the deraining task. (a) The paradigm of the most modern networks that incorporate both local and global residual connections. (b) The restored image output by the state-of-the-art model, Restormer (Zamir et al., 2022). It still exhibits visible traces of rain. (c) The restored image output by our LIR. It is cleaner and solves the problem of residual connections carrying degradation. (d) The visualization of intermediate feature maps of PReNet (Ren et al., 2019), SFNet (Cui et al., 2022), IRNeXt (Cui et al., 2023), Restormer, and LIR.
1. Introduction
@@
Image Restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, etc. In recent years, we have witnessed the great process of deep learning methods (Liang et al., 2021; Zamir et al., 2021; Tu et al., 2022; Luo et al., 2020; Poirier-Ginter and Lalonde, 2023; Wang et al., 2022; Zhang et al., 2023c) in IR. However, many existing works tend to focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task. This approach often leads to parameter redundancy and computations unnecessary. Instead, our goal is an efficient and lightweight network design for various IR tasks.
@@
To address the issue of vanishing gradient and stabilize training, as shown in Figure 1 (a), most networks usually incorporate local and global residual connections. However, these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output. To validate our idea, we conduct visualization experiments on deraining task using Restormer (Zamir et al., 2022), SFNet (Cui et al., 2022), IRNeXt (Cui et al., 2023), and PReNet (Ren et al., 2019). We select intermediate feature maps near the input and output of the networks for visualization because the semantics of the middle layer are too abstract. Results are shown in Figure 1 (d). Notably, it is evident that each feature map of these networks and the final output of the state-of-the-art model (Figure 1 (b)), Restormer, all contain severe degradation (rain). To overcome this issue, we use an ingenious structural design, using transpose branches in the LIR and LAA (Figure 2), that effectively eliminates the degradation in the local and global residual connections. The intermediate feature maps of our LIR (Figure 1 (d)) show better visual performance compared with the rest of the networks, and the restored image output by our LIR (Figure 1 (c)), which is more cleaner than the Restormer (Figure 1 (b)). More visual comparison of our LIR and these networks see Supplementary Materials.
@@
Moreover, we observe another issue in the visualization experiment. As shown Figure 1 (d)), the intermediate feature maps of the PReNet, SFNet, IRNetXt, and Resormer not only exhibit blurred contours but also contain a considerable amount of noise. The more important is that the high-frequency information has been lost. These observations indicate that although many state-of-the-art networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and missing to address the core issues of image restoration, thus resulting in poor visual quality. Based on this insight, we propose the Adaptive Filter (Figure 3) (a component of the LAA) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
@@
Transformer-based models have demonstrated impressive performance in various tasks, both at high-level (Dosovitskiy et al., 2020; Dong et al., 2022; Liu et al., 2021b; Touvron et al., 2021) and low-level (Chen et al., 2022b, 2021b; Wang et al., 2022; Tsai et al., 2022; Li et al., 2023). However, there exist different opinions. (Dong et al., 2021) argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs). In addition, (Liu et al., 2021a; Tolstikhin et al., 2021; Touvron et al., 2022) share a similar viewpoint that highlights the potential of MLPs beyond transformers. On the other hand, (Trockman and Kolter, 2022) suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy. Thus, excessive computations can be avoided. Furthermore, (Guo et al., 2023; Liu et al., 2022; Ding et al., 2022; Zhou et al., 2022) take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer. Satisfactory results are achieved and excessive computations are avoided. Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module (Figure 4) (an component of the Attention Block) that aims to capture global information of the features and keeps parameters and computations friendly.
@@
In this paper, we propose a Lightweight Image Restoration network called LIR (Figure 2) to efficiently remove degradation and adapt to different IR tasks. The key component of LIR is the novel Lightweight Adaptive Attention Block (LAA) consisting of several proposed Adaptive Filters and Attention Blocks. Where the Attention Block is formed by proposed Patch Attention module (Figure 4), ResCABlock, and ResBlock (details can be found in Figure 2). Thanks to the Adaptive Filter and Attention Block, the LAA Block is capable of adaptively sharpening contours, removing degradation, and capturing global information in various Image Restoration scenes in a computation-friendly manner. Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and LAA Block. This subtraction allows us to obtain the degradation, which is then expanded through the Transpose operation. After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection. Experiments demonstrate that our LIR achieves the state-of-the-art SSIM and comparable PSNR to state-of-the-art models on the deraining task, and comparable performance to state-of-the-art models with a parameter size of about 30% on the denoising, dehazing, and deblurring tasks.
@@
  • Through an ingenious structural design, the problem that local and global residual connections bring degradation is addressed, which is overlooked by modern networks.
  • Given the problem of the loss of high-frequency information observed in the visualization experiments, we propose the Adaptive Filter (Figure 3) as a solution to adaptively extract high-frequency information and enhance object contours.
  • To give LIR the ability to extract global information, we propose the Patch Attention module (Figure 4), which is more computation-friendly than the self-attention of the transformer.
@@
2. Related Work
@@
Figure 2. The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Lightweight Adaptive Attention (LAA) Blocks. Thanks to the Adaptive Filter, Patch Attention, ResCABlock, and the ResBlock, the LAA Block is endowed with the ability to extract a variety of information efficiently.
2.1. Image Restoration
In recent years, IR based on deep learning methods (Du et al., 2020; Zhang et al., 2023a; Xie et al., 2021; Li et al., 2020; Zhang et al., 2022; Huang et al., 2023) have shown promising results. The transformer architectures achieve adaptive spatial aggregation at the cost of heavy computation. To address this challenge, a Kernel Basis Attention (KBA) module has been proposed by KBNet (Zhang et al., 2023c) in the CNN. It adaptively aggregates spatial neighborhood information in a kernel-wise manner, thereby reducing computation. ADMS (Park et al., 2023) proposes a Filter Attribution method based on Integral Gradient (FAIG) (Xie et al., 2021) to identify the contributions of the filters to remove specific degradations. By masking filters with minimal impact on degradation removal, the network’s efficiency in removing degradations is improved. IDR (Zhang et al., 2023a) proposes a learnable Principal Component Analysis (PCA) and views various IR tasks as multi-task learning methods to obtain network priors and remove degradations. MPRNet (Zamir et al., 2021) introduces the Multi-Stage Progressive Restoration method, which is based on the U-Net architecture and follows a multi-input multi-output style. In addition, the contrastive learning paradigm is utilized by AirNet (Li et al., 2022a) to train the network to deal with different unknown degradations. Through the prompting of the contrastive learning paradigm, the latent features in the network gradually away from the degradation space and close to the clean space. The above works either involves a large number of parameters or a complex training process. Instead, our emphasis is on lightweight and efficient design while ensuring good performance.
2.2. Vision Transformer for Image Restoration
IPT (Chen et al., 2021b) is the first work that introduces ViT (Dosovitskiy et al., 2020) to low-level tasks to achieve image super-resolution, denoising, and deraining. The computational complexity of the self-attention part in ViT is $O(N^{2}D)$, where $N=H$$\times$$W$ represents the length of the patch sequence and $D$ represents the dimension of the sequence. Although vision transformers are capable of capturing long-range dependencies, the self-attention component imposes a significant computing burden. The computational complexity of the self-attention increases quadratically with the input size. To mitigate this issue, TransWeather (Valanarasu et al., 2022) introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$. Uformer (Wang et al., 2022) and StripFormer (Tsai et al., 2022) proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size. To further reduce the computational complexity, (Kong et al., 2023) estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL (Li et al., 2023) proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention. The computational complexities are reduced to $O(NlogND)$ and $O(NMD)$ respectively, where M represents the size of feature map. However, the computation required for self-attention remains unaffordable. To address this issue, we propose a novel Patch Attention module to approximate the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation for computational savings.
3. Method
@@
Figure 3. The detailed architecture of the proposed Adaptive Filter. It is a reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
Figure 4. The flow of the proposed Patch Attention.
3.1. Lightweight Image Restoration
The overall structure of LIR is illustrated in Figure 2, which primarily consists of Lightweight Adaptive Attention (LAA) Blocks. Both LIR and LAA Block adopt a down-up style to reduce computation and save parameters, that is downsizing the feature map in the beginning and upsizing at the end. To preserve low-frequency information and ensure training is stable, we add two global residual connections and one local residual connection into LIR. For the global connections, one connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv. For the local one, it connects the head and the tail.
As mentioned in Section 1, these residual connections can bring degradation. To address this, we subtract the output of the fourth LAA Block from the output of the first LAA Block to obtain the corresponding degradation. We then amplify the degradation through the Transpose operation and subtract it from the residual connection to obtain a clean connection. The same approach is used to remove degradation in the local residual connection within the LAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation. The reason we adopt this method is that the feature map of the shallow layer contains more degradation compared to a deep layer. Thus, this design allows the network to dynamically perceive this difference between shallow and deep layers during the training process and inference stage. Now, the problem that local and global residual connections bring degradation is addressed. Next, we illustrate how we can achieve lightweight and efficiency through the Adaptive Filter and Attention Block in detail.
3.2. Adaptive Filter
@@
To achieve lightweight and efficiency, we first introduce the Adaptive Filter (Figure 3), inspired by Edge-oriented Convolution Block (ECB) (Zhang et al., 2021b). The rich filters of different kinds allow the Adaptive Filter to enhance the object contour, remove degradation, and extract useful high-frequency information. Furthermore, the Adaptive Filter can be reparameterized as one $3\times 3$ and one high-pass filter after training to achieve lightweight.
Figure 5. The visual comparisons for LIR, Restormer (Zamir et al., 2022), SFNet (Cui et al., 2022), and HINet (Chen et al., 2021a). LIR produces state-of-the-art visual results.
@@
Our Adaptive Filter significantly outperforms ECB in three ways. Firstly, the ECB introduces more convolution operations during training leading to more training costs (i.e. parameters and computations). Instead, our Adaptive Filter only requires 10 training parameters resulting in smaller training costs. Secondly, the Adaptive Filter can be applied to various IR scenes and dynamically adjust the dependence on different filters by adjusting parameters $\alpha$. However, the ECB treats each branch equally and can not adapt to various IR scenes. Lastly, the Adaptive Filter offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
@@
Each filter in the Adaptive Filter has its own advantages and disadvantages. Thus, we use the learnable parameters $\alpha$ to adaptively maximize their advantages and minimize their disadvantages in different LAA blocks of the LIR. Specifically, in these filters, the Roberts filter is used to detect the pixel changes in the features by the differential method. It is effective in extracting prominent edges but fails to detect details. Where Roberts-x and Roberts-y represent the detection in the horizontal direction and vertical direction; The Sobel filter extracts edge information by approximating first-order differentials discretely. It has some resistance to noise and a good ability to detect details. However, the Sobel filter may cause edge thinning and false detection in scenes with a mass of noise; The Laplacian filter extracts edge information by approximating second-order differentials discretely. It can detect more detailed edges and textures than Sobel but is more sensitive to noise and may result in a bilateral effect on the edges. Where Laplacian-4 and Laplacian-8 represent 4 domains and 8 domains; The high-pass filter obtains high-frequency information by subtracting the features after Gaussian blur processing from the original features. This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the Adaptive Filter for stages where the features have scarce noise and a clear outline.
3.3. Attention Block